OpenStudy (anonymous):

find the limit (e^(2x)+2x)^(1/x) as x->infinity

OpenStudy (anonymous):

\[\text{Limit}\left[\left(e^{2 x}+2 x\right)^{\frac{1}{x}},x\to \text{Infinity}\right]=e^2 \]